# Dividing Indivisible Items for the Benefit of All: It is Hard to Be Fair Without Social Awareness

### v1

@@ 6 Discussion
One obvious direction for future research is to relax the SIM constraint, as someone can argue that SIM is too restrictive and this is the reason for non-existence of fair allocations for socially unaware agents; Flammini, Greco, and Varricchio (2025) show that we cannot get better than $\mathcal{O}\left({{n}}\right)$}-approximation for SIM under EF1 fairness. What if we replace SIM with a Pareto optimal allocation with respect to social impact? Unfortunately, there is no positive news for this: the proof of Theorem 3 shows that the problem remains NP-complete.
@@
A complementary approach to the question above is to further relax the fairness criterion. Does there always exist an allocation that is SIM and approximately fair, for some established fairness notion? A different idea is to try to strengthen the fairness guarantee for socially aware agents. Can we prove the existence of SIM and EFX allocations with socially aware agents for the settings that admit EFX allocations? What about SIM allocations satisfying epistemic EFX, that are guaranteed to exist and can be efficiently found (Aziz et al. 2018; Caragiannis et al. 2023; Akrami and Rathi 2025)?
@@
Another exciting avenue, which we believe deserves further study, is the setting where we have chores to allocate (Aziz 2016; Aziz et al. 2017, 2022). This is a very natural setting with many practical applications; think of the admin tasks in a CS department. Our initial results provided in the appendix show that, again, social unawareness does not allow for SIM and EF1 allocations. Besides this, even with social awareness, it is unclear whether we can always find SIM-EF1 allocations. In fact, the round-robin approach fails when we have chores.
Similarly, it is unclear whether algorithms based on envy-cycle elimination can be extended to SA-EF1 for chores. The main issue is that in the setting of chores, it is possible that agent $i$ does not SA-envy bundle ${A}_{j}$, when agent $j$ has the bundle, as agent $j$ has a smaller impact on the bundle of agent $i$; however, if ${A}_{j}$ is moved to some other agent $k$, who has same social impact for the $i$-th’s bundle as $i$, the agent $i$ can start SA-envying the same bundle. This is a problem, as obtaining SIM allocation requires eliminating the cycle only in the SA-envy graph. However, this means that it is possible that the agent to whom we allocate a chore already envies some bundle, just not SA-envies it on the current agent. Moving this bundle then can introduce envy that is not SA-EF1. Hence, the discussion above is a strong indication that if SA-EF1 allocations do indeed exist for chores, they require a different algorithmic technique.
Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002854). This work was co-funded by the European Union under the project Robotics and advanced industrial production (reg. no. CZ.02.01.01/00/22_008/0004590). AD acknowledges the support of the EPSRC grant EP/X039862/1. ŠS acknowledges the additional support of the Grant Agency of the Czech Technical University in Prague, grant No. SGS23/205/OHK3/3T/18.

### v2

@@ 6. Discussion
One obvious direction for future research is to relax the SIM constraint, as someone can argue that SIM is too restrictive and this is the reason for non-existence of fair allocations for socially unaware agents; (FlamminiGV2025, 29) show that we cannot get better than $\mathcal{O}\left({{n}}\right)$}-approximation for SIM under EF1 fairness. However, what if we replace SIM with a Pareto optimal allocation with respect to social impact? Unfortunately, there is no positive news for this: the proof of Theorem 4 shows that the problem remains NP-complete, even when there are only two agents.
@@
A complementary approach to the question above is to further relax the fairness criterion. Does there always exist an allocation that is SIM and approximately fair, for some established fairness notion? A different idea is to try to strengthen the fairness guarantee for socially aware agents. Can we prove the existence of SIM and EFX allocations with socially aware agents for the settings that admit EFX allocations in the “standard” model? What about SIM allocations satisfying epistemic EFX, that are guaranteed to exist and can be efficiently found (AzizBCGL2018, 6, 19, 2)?
@@
Another exciting avenue, which we believe deserves further study, is the setting where we have chores to allocate (Aziz2016, 5, 11, 7). This is a very natural setting with many practical applications; think of the admin tasks in a CS department. Our initial results provided in the appendix show that, again, social unawareness does not allow for SIM and EF1 allocations. Besides this, even with social awareness, it is unclear whether we can always find SIM-EF1 allocations. In fact, we have created examples where round-robin approach, that gives us SA-EF1 for goods by Theorem 3, fails when we have chores.
Example 0.
Consider an instance with two agents with identical valuation $v$ and five items. For the first two items, both agents have value $v(g_{1})=v(g_{2})=-100$ and $s_{i}(g_{1})=s_{i}(g_{2})=1$ for $i\in\{1,2\}$. For the remaining three items $v(g_{3})=v(g_{4})=v(g_{5})=-1$ but $s_{1}(g_{j})=0$ and $s_{2}(g_{j})=1$ for $j\in\{3,4,5\}$. Clearly, if we allocate the chores in round-robin fashion such that in each round agents pick their best chore for which they maximize the social impact, if such chore exists, then we end up with allocation $(\{g_{1},g_{2}\},\{g_{3},g_{4},g_{5}\})$. However, since $s_{1}(\{g_{1},g_{2}\})=s_{2}(\{g_{1},g_{2}\})$ and $v(\{g_{1}\})=v(\{g_{2}\})=-100<v(\{g_{3},g_{4},g_{5}\})=-3$, agent 1 envies agent 2.
Similarly, it is unclear whether algorithms based on envy-cycle elimination can be extended to SA-EF1 for chores. The main issue is that in the setting of chores, it is possible that agent $i$ does not SA-envy bundle ${A}_{j}$, when agent $j$ has the bundle, as agent $j$ has a smaller impact on the bundle of agent $i$; however, if ${A}_{j}$ is moved to some other agent $k$, who has same social impact for the $i$-th’s bundle as $i$, the agent $i$ can start SA-envying the same bundle. This is a problem, as obtaining SIM allocation requires eliminating the cycle only in the SA-envy graph. However, this means that it is possible that the agent to whom we allocate a chore already envies some bundle, just not SA-envies it on the current agent. Moving this bundle then can introduce envy that is not SA-EF1. Hence, the discussion above is a strong indication, at least to us, that if SA-EF1 allocations do indeed exist for chores, they require a different algorithmic technique in order to resolve it.
Acknowledgements.
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002854). This work was co-funded by the European Union under the project Robotics and advanced industrial production (reg. no. CZ.02.01.01/00/22_008/0004590). Argyrios Deligkas acknowledges the support of the EPSRC grant EP/X039862/1. Šimon Schierreich acknowledges the additional support of the Grant Agency of the Czech Technical University in Prague, grant No. SGS23/205/OHK3/3T/18. [Uncaptioned image]